on 4. C ompuspeed is an agency-operation selling a computer software package for a major computer company. The budget has beePlease, I need the answer and the approaches step by step( all the information about how to get the data) in an easy way for my understanding Thanks!
Answer #1

answer

a) contribution/sales ratio = 250000/750000 = 1:3

contribution = sales-variable cost

= 750000-500000 = 250000

b) break even point

  • in units = fixed cost / (sales per unit- variable cost per unit)

= 150000/(150-100)

= 3000 units

fixed cost = 70000 + 80000 = 150000

sales per unit = 750000/5000 = 150

variable cost per unit = 500000/5000 = 100

  • in sales = price x break even point in units

= 3000 x 150 = 450000

c) margin of safety

  • in sales = actual sales - breakeven sales

= 750000 - 450000 = 300000

  • in units = margin of safety in sales/price

= 300000/150 = 2000

d) new price = 150 add 5% = 157.5

new variable cost per unit = 100 add 10% = 110

new profit = 100000 add 50% = 150000

let units be x

profit = 157.5x - 110x - 150000 = 150000

x = 6316 approx

e) now units sold = 5000 + 6000 = 11000

profit = 50 x 11000 - 150000 - 25000 = 375000

break even units = fixed cost/contribution per unit

= 175000/50 = 3500 units